# Variance-Aware Noisy Training: Hardening DNNs against Unstable Analog Computations

### v1

@@ Abstract
This study underscores the necessity of Noisy Training while revealing its fundamental limitations in the presence of dynamic noise. To address these challenges, we propose Variance-Aware Noisy Training, a novel approach that mitigates performance degradation by incorporating noise schedules which emulate the evolving noise conditions encountered during inference. Our method substantially improves model robustness, without training overhead. Through experiments on image classification tasks in dynamic noise environments, we demonstrate a significant increase in robustness, from 72.3% with conventional Noisy Training to 97.3% with Variance-Aware Noisy Training on CIFAR-10 and from 38.5% to 89.9% on Tiny ImageNet.
@@ 1 Introduction
Deep neural networks (DNNs) have driven remarkable advancements in a wide array of machine learning applications, from computer vision and natural language, speech and signal processing. These breakthroughs are largely enabled by digital compute platforms, such as graphics processing units (GPUs) or specialized accelerators, which offer high throughput and flexibility. However, as DNNs grow in scale and are increasingly deployed in energy-constrained environments, the quest for more efficient hardware solutions becomes paramount. In addition, CMOS technology scaling is stuttering, thus alternative approaches to maintain performance scaling have to be found. Analog computing architectures replace discrete quantities with continuous ones, leveraging the inherent properties of physical systems to perform computations efficiently. These architectures enhance computational efficiency, reduce data movement overhead, and enable highly parallel multiply-accumulate (MAC) operations, while their most significant advantage lies in superior energy efficiency [8, 36].
@@ 3 Neural Networks and Noisy Environments
In order to establish the effectiveness of our proposed method, experiments are performed for different networks and datasets. For the initial and comprehensive evaluation, we perform image classification on CIFAR-10 [25], CINIC-10 [9] and Tiny ImageNet [27]. For CIFAR-10, we evaluate two model architectures: LeNet-5 [28] and ResNet-18 [19]. While for Tiny ImageNet and CINIC-10, LeNet-5 undersized and thus we focus on ResNet-18 and ResNet-50.
LeNet-5 and ResNet-18 use initial learning rates of 0.001 and 0.01 on CIFAR-10, respectively, while ResNet-18 and ResNet-50 use 0.001 on Tiny ImageNet and CINIC-10. All models are trained with Adam and cosine learning rate decay, using a batch size of 128 for 400 epochs.  [^1] The code is available at: https://anonymous.4open.science/r/VANT-8B7B
@@ 3.1 Global Noise Injection and Noisy Training
To simulate the noisy environment present in analog hardware, we inject noise at a global level during model computation. In this work, we follow the Walking Noise [3] methodology, which focuses on injecting noise at the activations. We consider additive Gaussian noise, due to the distributions widespread observation in natural processes and success in other works on noise injection [43]. To inject noise without bias, we sample with zero mean, i.e. $\mathcal{N}\left(0,\sigma\right)$, with $\sigma$ being the standard deviation of the noise. A schematic of the noise injection is shown in Figure 1. By varying the noise level $\sigma$ during inference, we assess the model’s robustness under noisy environments of different intensities.
@@ 3.2 Evaluation: Quantifying Robustness under Noise
$\text{rAUC}=\frac{\text{AUC of the method}}{\text{AUC of the upper bound curve}}[\%]$
@@ 4 Noisy Training: Strong but Not Flawless
As shown, training with noise injection significantly enhances robustness compared to a baseline model trained without noise. However, it is important to understand how other robustness-enhancing methods contribute to overall performance. These and the limits of Noisy Training will be explored in the following.
@@ 4.2 Limitations of Noisy Training
This observation leads to a crucial conclusion: Noisy Training is only effective when the noise characteristics are precisely known. If the noise level during training does not align with the actual noise encountered during deployment, the model’s robustness can be significantly compromised. This leads to the central research question of this work: {mdframed}[backgroundcolor=gray!10, linecolor=black, linewidth=1pt, roundcorner=5pt] How can we train models that remain robust across an entire fleet of devices, each potentially exhibiting different noise strengths over time?
@@ 5.1 Methodology: Variance-Aware Noisy Training
Here $\sigma_{\text{train}}$ is an extrinsic parameter, representing the known noise characteristic of a given hardware target. VANT additionally introduces $\alpha$ and $\theta$. Here $\alpha$ is utilized to correct for an offset introduced through sampling $\sigma_{\text{var}}$, which right shifts the accuracy curve. And $\theta$ adjusts Noisy Training to the time variations of a given accelerator. Looking forward to Sections 5.2 and 5.3, we note that VANT is rather insensitive to $\alpha$, while $\theta$ strongly depends on the chosen $\sigma_{\text{train}}$.
@@ 5.2 Experiments on CIFAR-10
Initially we evaluate $\alpha$ and $\theta$ individually to explore their general behavior (Fig. 4).
@@
For any setting of $\alpha$ and $\theta$, the robustness (rAUC) of VANT is significantly better compared to standard Noisy Training. And results are similar for LeNet-5. While variations in $\alpha$ appear to have little impact on the overall robustness as seen in Figure 4(b), $\theta$ plays a significant role in both the overall robustness and shape of the curve, see Figure 4(a). Furthermore, in Figure 4(a) it also becomes apparent that the method does not necessarily preserve the peak accuracy of standard Noisy Training at $\sigma_{\text{train}}$. This is caused by the variation of $\theta$, which shifts the point of optimal accuracy for VANT further to the right. Thus, during the evaluation of $\alpha$ and $\theta$ values this asymmetric influence needs to be considered. This is accomplished by measuring how close the accuracy at $\sigma_{\text{train}}$ is to standard Noisy Training, since ideally VANT should preserve all advantages of Noisy Training. In order to quantify this behavior, a new metric is introduced: Preserved Accuracy. It measures by how much the accuracy of VANT and standard Noisy Training differ at the noise injection point of $\sigma_{\text{train}}$. Ideally a setting of $\alpha$ and $\theta$ can be found, which keeps this metric at zero or higher, perfectly preserving the accuracy of standard Noisy Training.
@@
In the case of Figure 5 we select $\alpha=0.45$ and $\theta=0.25$ as the optimal parameters.
Since different analog accelerators require different initial noise levels, we now explore how VANT behaves for different $\sigma_{\text{train}}$. As VANT is largely invariant to $\alpha$ we fix it to $0.45$, as a middle ground for the sweet spot from Fig. 5(b).
@@
Similarly to Fig. 5, Fig. 6 shows both robustness and preserved accuracy. However, in this case the x-axis denotes the change in $\sigma_{\text{train}}$, e.g. different hardware accelerators, and the y-axis explores the behavior of $\theta$. Again, there is a trade-off to be made between the preserved accuracy and the robustness.
@@ 5.3 Generalizing to Complex Data: CINIC-10 & Tiny ImageNet
Results for both quality metrics under consideration are shown in Figure 7. Looking first at the preserved accuracy, one can observe that the accuracy is generally preserved across datasets and models. A notable exception is $\sigma=0.3$, where the preserved accuracy drops across all experiments and increases in robustness are also low. We postulate that for this specific setting $\theta$ was ill-chosen. Further investigating the robustness in Figure 7(a), we observe that VANT improves the robustness for all models, datasets and strengths of injected noise. However the effect is not consistent across the whole range of injected noise. Instead VANT really significantly improves the robustness for noise strengths of $\sigma_{\text{train}}\geq 0.5$. The underlying reason is that in order to achieve high accuracy under low noise, regions of higher noise are less prominently sampled. This results in subpar performance for much of the investigated range by the rAUC metric. Notably this behavior is largely inherited from Noisy Training.
For $\sigma_{\text{train}}\geq 0.5$ however, VANT is considerably more robust. Interestingly the largest improvements can be found on Tiny ImageNet with ResNet-50, the most complex dataset and model investigated. However also for ResNet-18 and CINIC-10 significant improvements are visible.
@@ 6 Summary
Empirical evaluations demonstrate the effectiveness of VANT in improving robustness across different noise regimes and dataset complexities. For instance, under typical analog noise conditions, VANT increases robustness from 72.3% to 97.3% on CIFAR-10. On the more challenging Tiny ImageNet dataset, VANT similarly yields significant gains, improving performance from 38.5% to 89.9%.

### v2

@@ Abstract
This study underscores the necessity of Noisy Training while revealing its fundamental limitations in the presence of dynamic noise. To address these challenges, we propose Variance-Aware Noisy Training, a novel approach that mitigates performance degradation by incorporating noise schedules which emulate the evolving noise conditions encountered during inference. Our method substantially improves model robustness, without training overhead. Through experiments on image classification tasks in dynamic noise environments, we demonstrate a significant increase in robustness, from 79.3% with conventional Noisy Training to 97.6% with Variance-Aware Noisy Training on CIFAR-10 and from 32.4% to 99.7% on Tiny ImageNet.
@@ 1 Introduction
Deep neural networks (DNNs) have driven remarkable advancements in a wide array of machine learning applications, from computer vision and natural language, speech and signal processing. These breakthroughs are largely enabled by digital compute platforms, such as graphics processing units (GPUs) or specialized accelerators, which offer high throughput and flexibility. However, as DNNs grow in scale and are increasingly deployed in energy-constrained environments, the quest for more efficient hardware solutions becomes paramount. In addition, Complementary Metal-Oxide-Semiconductor (CMOS) technology scaling is stuttering, thus alternative approaches to maintain performance scaling have to be found. Analog computing architectures replace discrete quantities with continuous ones, leveraging the inherent properties of physical systems to perform computations efficiently. These architectures enhance computational efficiency, reduce data movement overhead, and enable highly parallel multiply-accumulate (MAC) operations, while their most significant advantage lies in superior energy efficiency [8, 36].
@@ 3 Neural Networks and Noisy Environments
In order to establish the effectiveness of our proposed method, experiments are performed for different networks and datasets. For the initial and comprehensive evaluation, we perform image classification on CIFAR-10 [25], CINIC-10 [9] and Tiny ImageNet [27]. For CIFAR-10, we evaluate two model architectures: LeNet-5 [28] and ResNet-18 [19]. While for Tiny ImageNet and CINIC-10, LeNet-5 is undersized and thus we focus on ResNet-18 and ResNet-50.
LeNet-5 and ResNet-18 use initial learning rates of 0.001 and 0.01 on CIFAR-10, respectively, while both ResNet-18 and ResNet-50 use a learning rate of 0.001 on Tiny ImageNet and CINIC-10. All models are trained with Adam and cosine learning rate decay, using a batch size of 128 for 400 epochs.  [^1] The code is available at: https://github.com/HAWAIILAB/VANT
@@ 3.1 Global Noise Injection and Noisy Training
To simulate the noisy environment present in analog hardware, we inject noise at a global level during model computation. In this work, we follow the Walking Noise [3] methodology, which focuses on injecting noise at the activations. We consider additive Gaussian noise, due to its widespread occurrence in natural processes and its demonstrated effectiveness in previous works on noise injection [43]. To inject noise without bias, we sample with zero mean, i.e. $\mathcal{N}\left(0,\sigma\right)$, with $\sigma$ being the standard deviation of the noise. A schematic of the noise injection is shown in Figure 1. By varying the noise level $\sigma$ during inference, we assess the model’s robustness under noisy environments of different intensities.
@@ 3.2 Evaluation: Quantifying Robustness under Noise
[…]
@@ 4 Noisy Training: Strong but Not Flawless
As shown, training with noise injection significantly enhances robustness compared to a baseline model trained without noise. However, it is important to understand how other robustness-enhancing methods contribute to overall performance. In the following, we explore these methods as well as the limitations of Noisy Training.
@@ 4.2 Limitations of Noisy Training
This observation leads to a crucial conclusion: Noisy Training is only effective when the noise characteristics are precisely known. If the noise level during training does not align with the actual noise encountered during deployment, the model’s robustness can be significantly compromised. This leads to the central research question of this work: How can we train models that remain robust across an entire fleet of devices, each potentially exhibiting different noise strengths over time?
@@ 5.1 Methodology: Variance-Aware Noisy Training
Here $\sigma_{\text{train}}$ is an extrinsic parameter, representing the known noise characteristic of a given hardware target. VANT additionally introduces two parameters: $\theta$ adjusts Noisy Training to the time variations of a given accelerator, while $\alpha$ is a calibration parameter for $\sigma_{\text{train}}$. Looking forward to Sections 5.2 and 5.3, we note that VANT is rather insensitive to $\alpha$, while $\theta$ strongly depends on the chosen $\sigma_{\text{train}}$.
@@ 5.2 Experiments on CIFAR-10
Initially we evaluate $\alpha$ and $\theta$ individually to explore their general behavior (Figure 4).
@@
For any setting of $\alpha$ and $\theta$, the robustness (rAUC) of VANT is significantly better compared to standard Noisy Training. And results are similar for LeNet-5. While variations in $\alpha$ appear to have little impact on the overall robustness as seen in Figure 4(b), $\theta$ plays a significant role in both the overall robustness and shape of the curve, see Figure 4(a). Furthermore, in Figure 4(a) it also becomes apparent that the method does not necessarily preserve the peak accuracy of standard Noisy Training at $\sigma_{\text{train}}$. The variation of noise, $\theta$, effectively increases the maximum noise observed, which shifts the point of optimal accuracy for VANT. Thus, during the evaluation of $\alpha$ and $\theta$ this influence needs to be considered. This is accomplished by measuring how close the accuracy at $\sigma_{\text{train}}$ is to standard Noisy Training, since ideally VANT should preserve all advantages of Noisy Training. In order to quantify this behavior, a new metric is introduced: Preserved Accuracy. It measures by how much the accuracy of VANT and standard Noisy Training differ at the noise injection point of $\sigma_{\text{train}}$. Ideally a setting of $\alpha$ and $\theta$ can be found, which keeps this metric at zero or higher, perfectly preserving the accuracy of standard Noisy Training.
@@
In the case of Figure 5, we select $\alpha=0.45$ and $\theta=0.25$ as the optimal parameters.
Since different analog accelerators require different initial noise levels, we now explore how VANT behaves for different $\sigma_{\text{train}}$. As VANT is largely invariant to $\alpha$ we fix it to $0.45$, as a middle ground for the sweet spot from Figure 5(b).
@@
Similarly to Figure 5, Figure 6 shows both robustness and preserved accuracy. However, in this case the x-axis denotes the change in $\sigma_{\text{train}}$, e.g. different hardware accelerators, and the y-axis explores the behavior of $\theta$. Again, there is a trade-off to be made between the preserved accuracy and the robustness.
@@ 5.3 Generalizing to Complex Data: CINIC-10 & Tiny ImageNet
Results for both quality metrics under consideration are shown in Figure 7. Looking first at the preserved accuracy, one can observe that the accuracy is generally preserved across datasets and models. A notable exception is $\sigma=0.3$, where the preserved accuracy drops across all experiments and increases in robustness are also low. We postulate that for this specific setting $\theta$ was ill-chosen. Further investigating the robustness in Figure 7(a), we observe that VANT improves the robustness for all models, datasets and strengths of injected noise. However the effect is not consistent across the whole range of injected noise. VANT provides less significant robustness improvements for noise strengths of $\sigma_{\text{train}}<0.5$. The underlying reason is that in order to achieve high accuracy under low noise, regions of higher noise are less prominently sampled. This results in subpar performance for much of the investigated range by the rAUC metric. Notably this behavior is largely inherited from Noisy Training.
For $\sigma_{\text{train}}\geq 0.5$, however, VANT is considerably more robust. Interestingly the largest improvements can be found on Tiny ImageNet with ResNet-50, the most complex dataset and model investigated, where the rAUC increases dramatically from 32.4% to 99.7% at $\sigma_{\text{train}}=0.9$. Nonetheless, siginificant improvements are also visible for ResNet-18 and CINIC-10.
@@ 6 Summary
Empirical evaluations demonstrate the effectiveness of VANT in improving robustness across different noise regimes and dataset complexities. For instance, under typical analog noise conditions, VANT increases robustness from 79.3% to 97.6% on CIFAR-10. On the more challenging Tiny ImageNet dataset, VANT similarly yields significant gains, improving performance from 32.4% to 99.7%.